# Synthesizing Moving People with 3D Control

### v1

@@ 4.2 Quantitative Results
Baselines. We compare our approaches with past and state-of-the-art methods: DreamPose [14], DisCo [28] and ControlNet [34] (for pose accuracy comparisons) [^1] We utilize the open-source official code and models provided by the authors to implement these baselines. We use diffusers [27] for ControlNet and Openpose extraction, and Detectron2 for DensePose extraction for DisCO. Since  Chan et al. [8] can only work for animating a specific person, we don’t compare with it in this paper.. We set inference steps as 50 for all the approaches for fair comparisons.
Comparisons on Frame-wise Generation Quality. We compare  3DHM with other methods on 2K2K test dataset, which is composed of 50 unseen human videos, at $256\times 256$ resolution. For each human video, we take 30 frames that represent the different viewpoints of each unseen person. The angles range from $0^{\circ}$ to $360^{\circ}$, we take one frame every $12^{\circ}$ to better evaluate the prediction and generalization ability of each model. As for DisCO, we strictly follow their setting and extract OpenPose for inference. As for DreamPose, we extract DensePose for inference. We evaluate the results and calculate the average score over all frames of each video. We set the background as black for all approaches for fair comparisons. We report the average score overall of the same 50 videos and show the comparisons in Table 1. We observe that 3DHM outperforms all the baselines in different metrics.
@@
DreamPose 35.06 0.80 245.19 0.18 2.12e-04
DisCO 35.38 0.81 164.34 0.15 1.44e-04
Ours 36.18 0.86 154.75 0.12 9.88e-05
Table 1: Quantitative comparison on frame-wise generation quality : We compare our method with prior works on pose condition generation tasks and measure the generation quality of the samples.
Comparisons on Video-level Generation Quality. To verify the temporal consistency of 3DHM, we also report the results following the same test set and baseline implementation as in image-level evaluation. Unlike image-level comparisons, we concatenate every consecutive 16 frames to form a sample of each unseen person on challenging viewpoints. The angles range from $150^{\circ}$ to $195^{\circ}$, we take one frame every $3^{\circ}$ to better evaluate the prediction and generalization ability of each model. We report the average score overall of 50 videos and show the comparisons in Table 2. We observe that 3DHM, though trained and tested by per frame, still embrace significant advantage over prior approaches, indicating superior performance on preserving the temporal consistency with 3D control.
[…]
DreamPose 113.96 950.40
DisCO 83.91 629.18
Ours 55.40 422.38
Table 2: Quantitative comparison on video-level generation quality.
Comparisons on Pose Accuracy. To further evaluate the validity of our model, we estimate 3D poses from generated human videos from different approaches via a state-of-the-art 3D pose estimation model 4DHumans. We use the same dataset setting mentioned above and compare the extracted poses with 3D poses from the target videos. Following the same comparison settings with generation quality, we evaluate the results and calculate the average score over all frames of each video. Beyond DreamPose and DisCO, we also compare with ControlNet, which achieves the state-of-the-art in generating images with conditions, including openpose control. Since ControlNet does not input images, we input the same prompts as ours ‘a real human is acting’ and the corresponding openpose as conditions. We report the average score overall of 50 test videos and show the comparisons in Table 3. We could notice that 3DHM could synthesize moving people following the provided 3D poses with very high accuracy. At the same time, previous approaches might not achieve the same performance by directly predicting the pose-to-pixel mapping. We also notice that 3DHM could achieve superior results on both 2D metrics and 3D metrics, even if DisCO and ControlNet are controlled by Openpose and DreamPose is controlled by DensePose.
@@
Table 3: Quantitative comparison on pose accuracy.
@@ 4.3 Qualitative Results
Our work focuses on synthesizing moving people, primarily for clothing and the human body. With the aid of 3D assistance, our approach has the potential to produce human motion videos in various scenarios. We consider challenging 3D poses and motions from 3 sources: 3D human videos, random YouTube videos, and text input.
@@
Motions from Random YouTube Videos. We test our model on very different motions from randomly downloaded Youtube videos for an unseen human. We display the results in Figure 4(b).
Motions from Text Inputs. We test our model on motions from arbitrary text prompts. We randomly input an unseen human photo and motions from random text inputs via a widely used human motion generative model (MDM [25]). We display the results in Figure 4(c).
(a) 3DHM with a random human photo and a random 3D pose of various viewpoints. We show that even if the person’s photo is from a side angle, our stage 1 can help reconstruct the full texture map, which could be used to obtain full body information. Stage 2 can add texture information based on a given input.
(b) 3DHM with a random human photo and motions from random YouTube Videos. This example is from Gene Kelly’s dancing video.
(c) 3DHM with a random human photo and motions generated from text inputs by MDM, a Human Motion Diffusion Model [25].
Figure 5: Qualitative results on different viewpoints of the same pose; motions from random videos and motions from text input.
5 Analysis and Discussion
5.1 Ablation Study
To further verify the components of our methods, we train on training dataset and test on test datasets. We extract the 3D rendered pose from these 50 test video tracks. Same with the settings in quantitative comparison, we calculate the average scores of PSNR, SSIM, VGG, L1, LPIPS among all the generated frames and targeted original frames and report the results on both frame-wise metric (Table 4), video-level metric (Table 5) and pose accuracy (Table 6). We find that both texture map reconstruction and appearance latents are critical to the model performance. Also, we notice that directly adding SMPL parameters into the model during training may not bring improved performance considering all evaluation metrics. This is presumably due to the imprecision of SMPL parameters, which could provide contradictory information throughout the diffusion training process if they are not incorporated correctly.
[…]
Default 36.18 0.86 154.75 0.12 9.88e-05
w/o Texture map 35.00 0.78 237.42 0.20 2.35e-04
w/o Appearance Latents 36.07 0.86 167.58 0.12 1.03e-04
adding SMPL parameters 36.42 0.87 157.60 0.12 8.87e-05
Table 4: Ablation study of Rendering Diffusion. We compare the frame-wise generation quality under different settings. We notice both texturemap reconstruction and appearance latents are critical to the model performance.
[…]
Default 55.40 422.38
w/o Texture map 113.97 632.67
w/o Appearance Latents 93.21 715.51
adding SMPL parameters 72.35 579.90
Table 5: Ablation study of Rendering Diffusion. We compare the video-level generation quality under different settings. We notice that although adding SMPL parameters achieve better performance on frame-wise setting but may yield worse temporal consistency than default settings.
Method MPVPE $\downarrow$ PA-MPVPE $\downarrow$
Default 41.08 31.86
w/o Texture map 92.94 59.18
w/o Appearance Latents 41.99 32.82
adding SMPL parameters 39.16 29.67
Table 6: Ablation study of Rendering Diffusion. We compare the pose accuracy under different settings.
5.2 2D Control and 3D Control
@@ 5.1 Ablation Study
We also compare the results of the official model from DreamPose and DisCO on a random person on a random real human photo which ensures distinct data distribution. We display the qualitative results of various viewpoints in Figure 6. DreamPose, DisCO, and 3DHM are all initialize the U-Net model with the pre-trained weights of Stable Diffusion. We notice that 3DHM can generalize well to unseen real humans though it is only trained by limited 3D humans. Since DreamPose requires subject-specific finetuning of the UNet to achieve better results, it cannot directly generalize well on a random human photo. As for DisCO, though it has been trained with an effective human attribute pre-training on multiple public datasets for better generalizability to unseen humans, still fails to synthesize people without the target pose. We assume this is because 3DHM adds rigid 3D control to better correlate the appearance to the poses, and preserve the body shape. Training with OpenPose or DensePose cannot guarantee the mapping between textures and poses, which makes it hard for the models to generalize.
Figure 6: Qualitative comparison with other 2D control approaches on a random real human photo (a Korean actress). We apply various 3D poses or the same 3D pose from different viewpoints. It could be noticed that 2D poses may not be able to capture the folding motion, and details of the human body. We could notice that our approach 3DHM could bridge this gap with 3D control.
@@ 5.3 Limitations
As 3DHM generates the frames of the human motion videos independently, there is no guarantee of consistency in time. For example, the clothing light may change between consecutive frames. One possible solution is to train the model to predict multiple frames simultaneously. Another possible solution is to condition the generation process on previously generated frames via stochastic conditioning [1]. Additionally, since 3DHM is trained on a dataset of 2K people, not all the detailed textures can be reconstructed completely during inference (e.g. unique logos on the clothes). We hypothesize this could be alleviated by training with more human data.
@@ 6 Conclusion
In this paper, we propose 3DHM, a two-stage diffusion model-based framework that enables synthesizing moving people based on one random photo and target human poses. A notable aspect of our approach is that we employ a cutting-edge 3D pose estimation model to generate human motion data, allowing our model to be trained on arbitrary videos without necessitating ground truth labels. Our method is suitable for long-range motion generation, and can deal with arbitrary poses with superior performance over previous approaches.
@@ Acknowledgement
We thank the Machine Common Sense project and ONR MURI award number N00014-21-1-2801. We also thank Google’s TPU Research Cloud (TRC) for providing cloud TPUs. We thank Georgios Pavlakos, Shubham Goel, and Jane Wu for the constructive feedback and helpful discussions.
@@ Appendix A Dataset Analysis
Figures 6(a) and 6(b) present the clothing type statistics of the training data (2,205 humans) and test data (50 humans). We count people based on four clothing categories: skirted attire, suit, casual wear, and others. In some cases, the clothing belongs to skirted attire and suits or casual wear, we will count this as skirted attire. For each clothing category, we tally two styles: tight-fitting and loose-fitting.

### v2

@@ 5.2 Quantitative Results
Baselines. We compare our approaches with past and state-of-the-art methods: DreamPose [16], DisCo [32] and ControlNet [41] (for pose accuracy comparisons) [^1] We utilize the open-source official code and models provided by the authors to implement these baselines. We use diffusers [31] for ControlNet and Openpose extraction, and Detectron2 for DensePose extraction for MagicAnimate and DreamPose. Since  Chan et al. [7] can only work for animating a specific person, we don’t compare with it in this paper.. We set inference steps as 50 for all the approaches for fair comparisons.
Comparisons on Frame-wise Generation Quality. We compare  3DHM with other methods on 2K2K test dataset, which is composed of 50 unseen human videos, at $256\times 256$ resolution. For each human video, we take 30 frames that represent the different viewpoints of each unseen person. The angles range from $0^{\circ}$ to $360^{\circ}$, we take one frame every $12^{\circ}$ to better evaluate the prediction and generalization ability of each model. As for DisCO, we strictly follow their setting and extract OpenPose for inference. We extract DensePose for inference DreamPose and MagicAnimate. We evaluate the results and calculate the average score over all frames of each video. We set the background as black for all approaches for fair comparisons. We report the average score of the same 50 videos and show the comparisons in Table 1. We observe that 3DHM outperforms all the baselines in different metrics.
@@
DreamPose 35.06 0.80 245.19 0.18 2.12e-04 113.96 950.40
DisCO 35.38 0.81 164.34 0.15 1.44e-04 83.91 629.18
MagicAnimate 32.57 0.65 300.66 0.29 5.80E-04 140.45 900.70
Ours 36.18 0.86 154.75 0.12 9.88e-05 55.40 422.38
Table 1: Quantitative comparison on generation quality: We compare our method with prior works on pose condition generation tasks and measure the generation quality of the samples.
Comparisons on Video-level Generation Quality. To verify the temporal consistency of 3DHM, we also report the results following the same test set and baseline implementation as in image-level evaluation. Unlike image-level comparisons, we concatenate every consecutive 16 frames to form a sample of each unseen person on challenging viewpoints. The angles range from $150^{\circ}$ to $195^{\circ}$, we take one frame every $3^{\circ}$ to better evaluate the prediction and generalization ability of each model. We report the average score overall of 50 videos and show the comparisons in Table 1. We observe that 3DHM, though trained and tested by per frame, still embraces significant advantage over prior approaches, indicating superior performance on preserving the temporal consistency with 3D control.
Comparisons on Pose Accuracy. To further evaluate the validity of our model, we estimate 3D poses from generated human videos from different approaches via a state-of-the-art 3D pose estimation model 4DHumans. We use the same dataset setting mentioned above and compare the extracted poses with 3D poses from the target videos. Following the same comparison settings with generation quality, we evaluate the results and calculate the average score over all frames of each video. Beyond DreamPose and DisCO, we also compare with ControlNet, which achieves the state-of-the-art in generating images with conditions, including openpose control. Since ControlNet does not input images, we input the same prompts as ours ‘a real human is acting’ and the corresponding openpose as conditions. We report the average score overall of 50 test videos and show the comparisons in Table 2. We could notice that 3DHM could synthesize moving people following the provided 3D poses with very high accuracy. At the same time, previous approaches might not achieve the same performance by directly predicting the pose-to-pixel mapping. We also notice that 3DHM could achieve better results on both 2D metrics and 3D metrics, even if DisCO and ControlNet are controlled by Openpose and DreamPose is controlled by DensePose.
@@
Table 2: Quantitative comparison on pose accuracy. We measure the pose accuracy in the generated images given the conditioned pose as the ground truth. We can see that our model is very accurate in persevering the poses in generated images.
5.3 Ablation Study
To further verify the components of our methods, we train on training dataset and test on test datasets. We extract the 3D rendered pose from these 50 test video tracks. Same with the settings in quantitative comparison, we calculate the average scores among all the generated frames and targeted original frames and report the results on both frame-wise metric (PSNR, SSIM, FID, LPIPS, L1), video-level metric (FID-VID, FVD) and pose accuracy (MPVPE, PA-MPVPE) in Table 4. We find that both texture map reconstruction and appearance latents are critical to the model performance. Also, we notice that directly adding SMPL parameters into the model during training may not bring improved performance considering all evaluation metrics.
Running Cost. Here we outline the comparison of parameters and running time with other methods in Table 1 using a single GPU A100. We show the comparison in Table 3.
Method Time (second/frame) Parameter
DreamPose 22.0 1.0B
DisCO 5.0 2.0B
MagicAnimate 10.0 2.0B
Ours 3.2 2.0B
Table 3: Comparison of running cost. We compare inference time for different models, and we can see that our models is faster in comparison with our models.
6 Analysis and Discussion
[…]
Default 36.18 0.86 154.75 0.12 9.88e-05 55.40 422.38 41.08 31.86
w/o Texture map 35.00 0.78 237.42 0.20 2.35e-04 113.97 632.67 92.94 59.18
w/o Appearance Latents 36.07 0.86 167.58 0.12 1.03e-04 93.21 715.51 41.99 32.82
adding SMPL parameters 36.42 0.87 157.60 0.12 8.87e-05 72.35 579.90 39.16 29.67
Table 4: Ablation study of Rendering Diffusion. We compare the frame-wise generation quality, video-level generation quality and the pose accuracy under different settings. We notice both texturemap reconstruction and appearance latents are critical to the model performance. The results show that although adding SMPL parameters achieve better performance on frame-wise setting but may yield worse temporal consistency than default settings. Note: we use bold to represent the best result and underline to represent the second-best result.
@@ 6.1 Qualitative Results
Our work focuses on synthesizing moving people, primarily for clothing and the human body. With the aid of 3D assistance, our approach has the potential to produce human motion videos in various scenarios. We consider challenging 3D poses and motions from 2 sources: 3D human videos and random YouTube videos. We utilize our model, which has been scaled up for real-world domains.
@@
Motions from Random YouTube Videos. We test our model on very different motions from randomly downloaded YouTube videos for an unseen human. We display the results in Figure 6(b). The results show that 3DHM can efficiently animate any person using random motion resources, accurately following the 3D poses from challenging motion sources.
6.2 Qualitative Comparison
@@
We also compare the results of the official model from DreamPose, DisCO, and MagicAnimate on a random person on a random real human photo which ensures distinct data distribution. We display the qualitative results of various poses on real human photos in Figure 7. We notice that 3DHM can generalize well to unseen real humans though it is only trained by limited 3D humans. Since DreamPose requires subject-specific finetuning of the UNet to achieve better results, it cannot directly generalize well on a random human photo. As for DisCO, though it has been trained with an effective human attribute pre-training on multiple public datasets for better generalizability to unseen humans, still fails to synthesize people without the target pose. MagicAnimate uses 3D pose features (DensePose) which better controls the appearance of input images. However it always suffers from severe artifacts on DensePose segmentation maps, which severely ruins the pose accuracy and consistency. We assume this is because 3DHM adds rigid 3D control to better correlate the appearance to the poses, and preserve the body shape. Training with OpenPose or DensePose cannot guarantee the mapping between textures and poses, which makes it hard for the models to generalize.
Figure 7: Qualitative comparison with other state-of-the-art approaches on a real human photo.
@@ 6.3 Limitations
As 3DHM has been trained with limited data (around 2K synthetic humans and 1K real humans), it might struggle to predict the texture details of the unseen side of the input human photo. However, we believe this issue can be mitigated by scaling up with more human data.
@@ 7 Conclusion
In this paper, we propose 3DHM, a two-stage diffusion model-based framework that enables synthesizing moving people based on one random photo and target sequence of human poses. A notable aspect of our approach is that we employ a cutting-edge 3D pose estimation model to generate human motion data, allowing our model to be trained on arbitrary videos without necessitating ground truth labels. Our method is suitable for long-range motion generation, and can deal with arbitrary poses with superior performance over previous approaches, by preserving the poses of the target motion, and clothing, face identities and smoother motion between frames.
@@ Acknowledgement
We thank the Machine Common Sense project and ONR MURI award number N00014-21-1-2801. We also thank Google’s TPU Research Cloud (TRC) for providing cloud TPUs. We thank Georgios Pavlakos, Shubham Goel, Jane Wu and Karttikeya Mangalam for the constructive feedback and helpful discussions.
@@ Appendix A Dataset Analysis
Figures 8(a) and 8(b) present the clothing type statistics of the synthetic training data (2,205 humans) and test data (50 humans). We count people based on four clothing categories: skirted attire, suit, casual wear, and others. In some cases, the clothing belongs to skirted attire and suits or casual wear, we will count this as skirted attire. For each clothing category, we tally two styles: tight-fitting and loose-fitting.